OpenStudy (anonymous):

Evaluate the integral ∫R2xydA, where R is the triangle 2x+y≤2, x≥0, y≥0

OpenStudy (jamesj):

Did you draw the region as recommended to you 12 hours ago?

OpenStudy (jamesj):

What then are the limits of integration?

OpenStudy (anonymous):

lol..i posed that last night..then had to leave

OpenStudy (anonymous):

so yea....i get the limits from 0 to 2, and 0 to 2x if that's correct?

OpenStudy (mr.math):

@james: lol, you remember the problems you solved last night?!

OpenStudy (jamesj):

No. Look carefully at the equation of that line. It's not y = 2x

OpenStudy (jamesj):

@M: it would seem so.

OpenStudy (anonymous):

oh it's 2x + y....so then x = 2/y?

OpenStudy (jamesj):

;-)

OpenStudy (jamesj):

Nooo. Draw the region. What are the three straight lines that form the boundary of the region R.

OpenStudy (anonymous):

okay hang on lemme look

OpenStudy (jamesj):

R triangle where 2x+y≤2, x≥0, y≥0

OpenStudy (anonymous):

okay..so if i'm reading this correctly...just elaborating...x is greater than 0...y is greater than zero putting this in quadrant 1....2x+y has to be less than 2?

OpenStudy (jamesj):

Yes, hence ....

OpenStudy (jamesj):

the three boundary lines of the triangle are ... what? And what are the coordinates of the vertices?

OpenStudy (anonymous):

how would i find that?

OpenStudy (anonymous):

i mean...it's probably simple and i'm overthinking it...but i don't see it

OpenStudy (mr.math):

Try to draw the line 2y+x=2.

OpenStudy (jamesj):

What does the straight line 2x + y = 2 look like?

OpenStudy (anonymous):

just a straight diagonal line

OpenStudy (jamesj):

Crossing the axes where? R = { (x,y) | 2x+y≤2, x≥0, y≥0 }

OpenStudy (mr.math):

Then focus on the region "below" this line, to the top of the x-axis, and to the right of the y-axis. That's what the given boundaries mean.

OpenStudy (anonymous):

at 2 and 1

OpenStudy (jamesj):

At (1,0) and (0,2)

OpenStudy (jamesj):

you've got to be more precise. Now, where is the third vertex/corner of the triangle?

OpenStudy (anonymous):

origin?

OpenStudy (jamesj):

Yes. That's where the lines x = 0 and y = 0 intersect.

OpenStudy (jamesj):

Given all that then, what are the limits of integration. You can do either y first then x, or vice versa. They give different answers. Be clear when you give your answer which order you are using.

OpenStudy (jamesj):

Actually, before you do that, draw yourself a diagram so it is clear to you what R looks like.

OpenStudy (anonymous):

i graphed it on my calculator

OpenStudy (anonymous):

and have it in front of me

OpenStudy (jamesj):

Ok. I recommend integrating y first; it's more intuitive in way. So y goes from what to what?

OpenStudy (anonymous):

0 to 2

OpenStudy (jamesj):

No. It does when x = 0 but not anywhere else.

OpenStudy (jamesj):

y = 0 to (some function of x)

OpenStudy (anonymous):

2x-y?

OpenStudy (jamesj):

y can't appear as a limit of a y integration. It's just a function of x. What is the equation of the straight line, y = f(x) = .... what?

OpenStudy (anonymous):

mx+b

OpenStudy (jamesj):

and what are the values of m and b? R = { (x,y) | 2x+y≤2, x≥0, y≥0 }

OpenStudy (anonymous):

2x-2

OpenStudy (jamesj):

can't be right. The line has negative slope.

OpenStudy (jamesj):

The equation of the line is 2x + y = 2. hence y = ...

OpenStudy (anonymous):

-2x+2

OpenStudy (jamesj):

Yes. Hence the limit of integration in the y-variable is y = 0 to (-2x + 2) Now, what's the limits of integration in the x-variable?

OpenStudy (jamesj):

This should have no variables; just numbers.

OpenStudy (jamesj):

x = .... to ....

OpenStudy (jamesj):

look at your diagram

OpenStudy (anonymous):

-1

OpenStudy (jamesj):

What? x = -1 is completely outside the region R.

OpenStudy (anonymous):

x is from 0 to 2

OpenStudy (anonymous):

sorry this server is getting laggy

OpenStudy (jamesj):

No. x only goes from 0 to 1.

OpenStudy (anonymous):

oh right...and y goes from 0 to 2

OpenStudy (jamesj):

Yes. Notice that if you had integrated over y = 0 to 2 and x = 0 to 1, that would be a rectangle, not a triangular region.

OpenStudy (jamesj):

This is why the first variable of integration in this case has to pick up on that triangular boundary.

OpenStudy (jamesj):

Now, given all that, your integral is \[ \int_0^1 \int_0^{-2x +2} 2xy \ dy \ dx \]

OpenStudy (jamesj):

Now integrate carefully and you'll have your answer.

OpenStudy (jamesj):

As an exercise, after you've found the answer, change the order of integration and make sure you get the same answer. The limits of integration will be x = 0 to (some function of y) y = 0 to 2

OpenStudy (anonymous):

i'm getting errors on my bounds

OpenStudy (anonymous):

i have to plug a b c and d for limits of integration

OpenStudy (jamesj):

What do you mean? In your calculator?

OpenStudy (anonymous):

no..i have to submit my answers online

OpenStudy (jamesj):

Well, I don't know what's up with your system, but the limits we have are definitely correct.

OpenStudy (jamesj):

...or I should say definitely correct given the order in which we're integrating. What order to they ask for in the question?

OpenStudy (mr.math):

sonofa posted the question again in a new post and I helped her out, I didn't know where exactly you've stopped.

OpenStudy (jamesj):

let me find it